if x+y=1, and x-y=1, then y = ?
Answer by ankor@dixie-net.com(22740)

Obviously y has to = 0, however prove it with elimination
:
x + y = 1
x - y = 1,
-------------Subtraction eliminates x and gives us y
0 + 2y = 0
y = 0